# Exploring The Missing Semantics In Event Modality

### v1

@@ III-C Semantic-E2VID
Cross-Modal Semantic Learning and Aligning. The goal of CFA is to explore semantic information by learning valuable knowledge with rich semantics from SAM. However, a natural modality mismatches as opposed to direct distillation. That is: the event feature $\mathcal{F}_{e}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$ encoded by the event encoder represent completely different information compared to the encoded frame feature with rich semantics $\mathcal{F}_{sam}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$ by SAM [^1] Note: SAM’s feature is extracted from the ground truth grayscale frames of the training dataset by SAM’s image encoder and was interpolated to the corresponding spatial dimension. No extra computational cost is needed during the inference stage. This also holds for the feature distillation learning process.. Therefore, directly applying the distillation loss between $\mathcal{F}_{e}$ and $\mathcal{F}_{\text{sam}}$ will cause a shifting in the event encoder learning and decoding process, which results in poor reconstruction, as shown by the ablation study in later sections. Thus, CFA is proposed to bridge the misalignment gap by mapping the encoded event feature $\mathcal{F}_{e}$ to SAM’s semantics without interfering with the event encoder. Specifically, as shown in Fig. 2 orange box, CFA takes the encoded event feature $\mathcal{F}_{e}$ as input and applies two 2D convolution operations on event features, combined with the GELU activation function, to map local event features to local frame features from SAM with rich semantics, and enhance semantic representation capabilities. Subsequently, Instance Norm is utilized to normalize features and align them with the statistical distribution of semantic features, reducing the impact of distribution discrepancies. Finally, to ensure the model can effectively utilize semantic information while preventing the propagation of error information, we employ distillation loss, which will be explained in a later section, to constrain the learned semantic features $\mathcal{F}_{\text{semantic}}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$ to approximate the ideal semantic features $\mathcal{F}_{\text{sam}}$ provided by SAM while avoiding misleading event encoder learning.
@@
Semantic-Aware Feature Fusion. While CFA enables Semantic-E2VID to explore rich semantic information, the next challenge is to fuse the learned semantics $\mathcal{F}_{\text{semantic}}$ with the encoded event feature $\mathcal{F}_{e}$ to generate encoded event features with rich semantics, readily for decoding into accurate video. To conquer this, SFF is employed, where its primary function is to enable deep interaction between $\mathcal{F}_{\text{semantic}}$ and $\mathcal{F}_{e}$ through a spatial attention mechanism and an adaptive fusion strategy, ensuring that semantic information exerts its influence at appropriate spatial locations. Specifically, as shown in Fig. 2 blue box, SFF first concatenates $\mathcal{F}_{\text{semantic}}$ from CFA with the original event features $\mathcal{F}_{e}$. Subsequently, 2D convolution, Batch Normalization, and GELU activation functions are employed to transform the features into an attention-like weight map. To ensure that semantic information acts at the correct position, we employ a self-attention mechanism [38] where global context information of the entire feature map is extracted via global average pooling, followed by actions such as 2D convolution and sigmoid activation functions, respectively. Finally, based on the attention weights, element-wise weighting is performed, and the adaptively generated semantic event features $\mathcal{F}_{\text{e}}^{\prime}$ is fed to later decoders for image reconstruction.
III-D Event To Video Supervision
@@
Semantic Perceptual E2V Supervision. The traditional perceptual reconstruction loss passes the reconstructed image and reference images through a pre-trained deep neural network trained for visual recognition tasks and uses the distance between the multiple layers of deep features from the network as a measure of the perceptual difference between the two images. However, the perception was done on the entire reconstructed frames, which may lead the E2V model to focus excessively on larger background regions and ignore small but critical objects with key semantics. To address this challenge, we utilize segmentation category results from SAM by employing its mask generation function [^2] https://github.com/facebookresearch/segment-anything to produce fine-grained semantic groupings. Specifically, single-point prompts are sampled in a grid across the image, from which SAM predicts multiple candidate masks. These masks are subsequently filtered for quality and deduplicated using non-maximum suppression [17]. Finally, we select the top $N$ non-background categories and resize them to construct a binary categorical mask $\mathcal{M}\in\mathbb{R}^{h\times w\times N}$. As illustrated in Fig. 3, to better supervise the key regions with rich semantics, we dot multiply the first, largest layer feature of the LPIPS features with $\mathcal{M}$ ($h=w=27$ to match the spatial size), to derive a masked LPIPS semantic feature layer. The remaining layers of the LPIPS feature remain unmasked to maintain the reconstruction quality of the background regions. Following previous E2V works [11, 12], we use the AlexNet [39] as the network for perceptual image patch similarity (LPIPS) [40], and we use the value $10$ for $N$ for all experiments. The Semantic Perception Supervision $\mathcal{L}_{\text{Semantic}}$ is calculated as below:
@@
Learning Semantics via Distillation. We adopt relational knowledge distillation similar to previous distillation works [33, 41] where the learned semantics $\mathcal{F}_{\text{semantic}}$ and semantics from SAM $\mathcal{F}_{\text{sam}}$ are flattened by their spatial dimension into feature maps $\mathcal{V}_{semantic}\in\mathbb{R}^{H\times W,C}$ and $\mathcal{V}_{sam}\in\mathbb{R}^{H\times W,C}$, after normalizing each feature map along the channel dimension. The spatial similarity matrix of students or teachers is calculated by inner product and the Distillation Loss can be calculated by the $L_{1}$ loss between the similarity matrices, as shown below:
@@
Ensuring Temporal Consistency. Following previous E2V works [12, 11], we deploy temporal consistency loss between the reconstructed images in adjacent time steps of the network. Specifically, temporal consistency loss is achieved by warping the previous reconstructed image using a ground truth optical flow to align with the current reconstructed image, and using the masked distance between these aligned images as a weight for temporal consistency, where the warping error between the previous true intensity image and the current true intensity image calculates the masked distance. The temporal consistency loss is calculated as:
@@
Final Optimization. Combining the three losses, our final optimization objective is:
@@
TABLE I: Quantitative comparison on ECD, MVSEC and HQF datasets. Lower is better ($\downarrow$) for MSE and LPIPS. Higher is better ($\uparrow$) for SSIM. The best and second-best indexes are marked in bold and underlined.
@@
IV Experiments
IV-A Training Protocols
Dataset. We randomly sampled images from the MSCOCO dataset [44] and utilized multi-object 2D renderer option of ESIM [45] to generate synthetic training datasets following [10]. These images serve as the background and foreground objects for the dataset. The dataset comprises $280$ sequences, each lasting $10$ seconds. Within the sequences, foreground images exhibit varying speeds and trajectories. Each sequence includes the generated event stream, ground truth images, and optical flow maps with an average rate of 51 Hz. Both the event camera and the frame camera have a resolution of $256\times 256$ pixels, while the contrast threshold range for event generation is $0.1$ to $1.5$.
Implementation Details. We utilized sequences of length $40$ to train our network, initializing the network parameters using He initialization. We set the batch size to 4 and trained the model for 400 epochs using the Adam optimizer, with a learning rate of $1\times e^{-5}$. We use $\lambda=1.8$, and $\alpha=50$ following [11, 12, 46] as hyper-parameters for the loss function. All networks were implemented in PyTorch and executed on a system equipped with a single RTX $3090$ GPU.
IV-B Evaluation Protocols
Evaluation on Data with Referenced Frames. Table I reports the quantitative results obtained using the aforementioned evaluation metrics. Our method consistently achieves state-of-the-art performance across most metrics. On the ECD and HQF datasets, it significantly outperforms HyperE2VID and EventMamba. For the MVSEC dataset, our approach attains the lowest MSE and highest SSIM scores, while also achieving competitive LPIPS results compared to transformer-based ET-Net with heavy computation cost [11]. These findings highlight the effectiveness of Semantic-E2VID in reconstructing high-fidelity images. By incorporating semantic priors during both feature encoding and decoding, our model can more accurately reconstruct structural and contextual information that corresponds to meaningful visual semantics in the scene.
Evaluation on Data with no Referenced GT Frames. Table II presents the quantitative results obtained from evaluating the methods on sequences from the HDR and NIGHT datasets. These benchmarks showcase the reconstruction ability under challenging conditions such as low light, overexposure, and extremely fast motions, where ground truth frames are not reliable. Our method achieves the best NIQE score in both benchmarks with challenging scenarios involving fast camera motion. This robustness stems from the learned semantic guidance, which helps maintain coherent object appearance and scene layout even when low-level signal quality is degraded.
TABLE II: NIQE metric ($\downarrow$) on HDR and NIGHT datasets. Best and second-best are in bold and underline.
@@ IV-B Evaluation Protocols
Evaluation on Color Frame Reconstruction. To further demonstrate the ability of our method for color image reconstruction, we use the CED dataset, which utilizes the DAVIS 346c camera to capture color frames and events. This dataset is essential for demonstrating the color reconstruction capability of the proposed method. As shown in Fig. 5, Semantic-E2VID produces more visually compelling reconstructions, particularly in scenes containing colored objects and fast-moving dynamics. The reconstructed images of Semantic-E2VID have higher contrast and clear edges, which proves the effectiveness and robustness of our proposed method.
@@
Qualitative Comparison. Figure 4 showcases qualitative reconstruction results produced by our Semantic-E2VID alongside all other SOTA methods on samples from datasets with referenced GT frames. It can be observed that methods such as E2VID, FireNet+, and SPADE-E2VID struggle with brightness consistency, leading to noticeably degraded visual quality. Our Semantic-E2VID demonstrates superior performance in preserving fine-grained object boundaries. This improvement is attributed to the incorporation of rich semantic information, which enhances the overall fidelity of the reconstruction. Furthermore, the outputs of Semantic-E2VID show a closer resemblance to the ground truth frames, delivering high visual quality while effectively mitigating reconstruction artifacts. Semantic-aware guidance allows our model to reconstruct high-level scene understanding, including object shapes and spatial relationships, yielding more realistic and interpretable video frames.
IV-C Qualitative Analysis
@@
IV-C1 Qualitative Analysis on Reconstruction Quality of Semantic Regions
To further demonstrate the effectiveness of our method in reconstructing semantically meaningful regions, we provide qualitative comparisons with state-of-the-art methods, including E2VID [12], E2VID+ [10], FireNet [9], FireNet+ [10], and ours. As shown in Figure 6, our method produces more accurate and sharper reconstructions in regions with rich semantic content (e.g., pedestrians, vehicles, or structured indoor objects).
@@ IV-C1 Qualitative Analysis on Reconstruction Quality of Semantic Regions
IV-C2 Qualitative Analysis on Downstream Segmentation Task
@@
IV-C3 Qualitative Analysis on Semantic Feature Learning
To better understand how our proposed semantic-guided learning module benefits the reconstruction process, we visualize the intermediate feature maps, $\mathcal{F}_{e}^{\prime}$, and final reconstructions from our method, and compare the visualization of the intermediate feature maps $\mathcal{F}_{e}$, and final reconstructions from the baseline model in Figure 8. We apply colormaps to visualize the feature activations, where warmer colors (e.g., red and yellow) indicate higher activation intensities, and cooler colors (e.g., blue) represent lower activations.

### v2

@@ 3.4 Semantic-E2VID
The goal of CFA is to explore semantic information by learning valuable knowledge with rich semantics from SAM. However, a natural modality mismatches as opposed to direct distillation. That is: the event feature $\mathcal{F}_{e}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$ encoded by the event encoder represent completely different information compared to the encoded frame feature with rich semantics $\mathcal{F}_{sam}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$ by SAM [^1] Note: SAM’s feature is extracted from the ground truth grayscale frames of the training dataset by SAM’s image encoder and was interpolated to the corresponding spatial dimension. No extra computational cost is needed during the inference stage. This also holds for the feature distillation learning process.. Therefore, directly applying the distillation loss between $\mathcal{F}_{e}$ and $\mathcal{F}_{\text{sam}}$ will cause a shifting in the event encoder learning and decoding process, which results in poor reconstruction, as shown by the ablation study in later sections. Thus, CFA is proposed to bridge the misalignment gap by mapping the encoded event feature $\mathcal{F}_{e}$ to SAM’s semantics without interfering with the event encoder. Specifically, as shown in Fig. 2 orange box, CFA takes the encoded event feature $\mathcal{F}_{e}$ as input and applies two 2D convolution operations on event features, combined with the GELU activation function, to map local event features to local frame features from SAM with rich semantics, and enhance semantic representation capabilities. Subsequently, Instance Norm is utilized to normalize features and align them with the statistical distribution of semantic features, reducing the impact of distribution discrepancies. Finally, to ensure the model can effectively utilize semantic information while preventing the propagation of error information, we employ distillation loss, which will be explained in a later section, to constrain the learned semantic features $\mathcal{F}_{\text{semantic}}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$ to approximate the ideal semantic features $\mathcal{F}_{\text{sam}}$ provided by SAM while avoiding misleading event encoder learning.
3.4.2 Semantic Feature Fusion.
@@
While CFA enables Semantic-E2VID to explore rich semantic information, the next challenge is to fuse the learned semantics $\mathcal{F}_{\text{semantic}}$ with the encoded event feature $\mathcal{F}_{e}$ to generate encoded event features with rich semantics, ready for decoding into an accurate video. To conquer this, SFF is employed, where its primary function is to enable deep interaction between $\mathcal{F}_{\text{semantic}}$ and $\mathcal{F}_{e}$ through a spatial attention mechanism and an adaptive fusion strategy, ensuring that semantic information exerts its influence at appropriate spatial locations. Specifically, as shown in Fig. 2 blue box, SFF first concatenates $\mathcal{F}_{\text{semantic}}$ from CFA with the original event features $\mathcal{F}_{e}$. Subsequently, 2D convolution, Batch Normalization, and GELU activation functions are employed to transform the features into an attention-like weight map. To ensure that semantic information acts at the correct position, we employ a self-attention mechanism [39] where global context information of the entire feature map is extracted via global average pooling, followed by actions such as 2D convolution and sigmoid activation functions, respectively. Finally, based on the attention weights, element-wise weighting is performed, and the adaptively generated semantic event features $\mathcal{F}_{\text{e}}^{\prime}$ is fed to later decoders for image reconstruction.
3.5 Semantic-Aware Supervision Event To Video Supervision
@@
3.5.1 Semantic Perceptual E2V Supervision.
The traditional perceptual reconstruction loss passes the reconstructed image and reference images through a pre-trained deep neural network trained for visual recognition tasks and uses the distance between the multiple layers of deep features from the network as a measure of the perceptual difference between the two images. However, the perception was done on the entire reconstructed frames, which may lead the E2V model to focus excessively on larger background regions and ignore small but critical objects with key semantics. To address this challenge, we utilize segmentation category results from SAM by employing its mask generation function [^2] https://github.com/facebookresearch/segment-anything to produce fine-grained semantic groupings. Specifically, single-point prompts are sampled in a grid across the image, from which SAM predicts multiple candidate masks. These masks are subsequently filtered for quality and deduplicated using non-maximum suppression [13]. Finally, we select the top $N$ non-background categories and resize them to construct a binary categorical mask $\mathcal{M}\in\mathbb{R}^{h\times w\times N}$. As illustrated in Fig. 3, to better supervise the key regions with rich semantics, we dot multiply the first, largest layer feature of the LPIPS features with $\mathcal{M}$ ($h=w=27$ to match the spatial size), to derive a masked LPIPS semantic feature layer. The remaining layers of the LPIPS feature remain unmasked to maintain the reconstruction quality of the background regions. Following previous E2V works [8, 32], we use the AlexNet [14] as the network for perceptual image patch similarity (LPIPS) [51], and we use the value $10$ for $N$ for all experiments. The Semantic Perception Supervision $\mathcal{L}_{\text{Semantic}}$ is calculated as below:
@@
3.5.2 Learning Semantics via Distillation.
We adopt relational knowledge distillation similar to previous distillation works [46, 29] where the learned semantics $\mathcal{F}_{\text{semantic}}$ and semantics from SAM $\mathcal{F}_{\text{sam}}$ are flattened by their spatial dimension into feature maps $\mathcal{V}_{semantic}\in\mathbb{R}^{H\times W,C}$ and $\mathcal{V}_{sam}\in\mathbb{R}^{H\times W,C}$, after normalizing each feature map along the channel dimension. The spatial similarity matrix of students or teachers is calculated by inner product and the Distillation Loss can be calculated by the $L_{1}$ loss between the similarity matrices, as shown below:
@@
3.5.3 Ensuring Temporal Consistency.
Following previous E2V works [32, 8], we deploy temporal consistency loss between the reconstructed images in adjacent time steps of the network. Specifically, temporal consistency loss is achieved by warping the previous reconstructed image using a ground truth optical flow to align with the current reconstructed image, and using the masked distance between these aligned images as a weight for temporal consistency, where the warping error between the previous true intensity image and the current true intensity image calculates the masked distance. The temporal consistency loss is calculated as:
@@
3.5.4 Final Optimization.
Combining the three losses, our final optimization objective is:
@@
4 Experiments
4.1 Training Protocols
4.1.1 Training Dataset.
We randomly sampled images from the MSCOCO dataset [19] and utilized multi-object 2D renderer option of ESIM [30] to generate synthetic training datasets following [36]. These images serve as the background and foreground objects for the dataset. The dataset comprises $280$ sequences, each lasting $10$ seconds. Within the sequences, foreground images exhibit varying speeds and trajectories. Each sequence includes the generated event stream, ground truth images, and optical flow maps with an average rate of 51 Hz. Both the event camera and the frame camera have a resolution of $256\times 256$ pixels, while the contrast threshold range for event generation is $0.1$ to $1.5$.
4.1.2 Implementation Details.
We utilized sequences of length $40$ to train our network, initializing the network parameters using He initialization. We set the batch size to 4 and trained the model for 400 epochs using the Adam optimizer, with a learning rate of $1\times e^{-5}$. We use $\lambda=1.8$, and $\alpha=50$ following [8, 32, 15] as hyper-parameters for the loss function. All networks were implemented in PyTorch and executed on a system equipped with a single RTX $3090$ GPU.
4.2 Evaluation Protocols
4.2.1 Evaluation on Data with Referenced Frames.
Table 1 reports the quantitative results obtained using the aforementioned evaluation metrics. Our method consistently achieves state-of-the-art performance across most metrics. On the ECD and HQF datasets, it significantly outperforms HyperE2VID and EventMamba. For the MVSEC dataset, our approach attains the lowest MSE and highest SSIM scores, while also achieving competitive LPIPS results compared to transformer-based ET-Net with heavy computation cost [8]. These findings highlight the effectiveness of Semantic-E2VID in reconstructing high-fidelity images. By incorporating semantic priors during both feature encoding and decoding, our model can more accurately reconstruct structural and contextual information that corresponds to meaningful visual semantics in the scene.
Table 1: Quantitative comparison on ECD, MVSEC and HQF datasets. Lower is better ($\downarrow$) for MSE and LPIPS. Higher is better ($\uparrow$) for SSIM. The best and second-best indexes are marked in bold and underlined.
@@ 4.2 Evaluation Protocols
4.2.2 Evaluation on Data with no Referenced GT Frames.
Table 2 presents the quantitative results obtained from evaluating the methods on sequences from the HDR and NIGHT datasets. These benchmarks showcase the reconstruction ability under challenging conditions such as low light, overexposure, and extremely fast motions, where ground truth frames are not reliable. Our method achieves the best NIQE score in both benchmarks with challenging scenarios involving fast camera motion. This robustness stems from the learned semantic guidance, which helps maintain coherent object appearance and scene layout even when low-level signal quality is degraded.
Table 2: NIQE metric ($\downarrow$) on HDR and NIGHT datasets. Best and second-best are in bold and underline.
@@
4.2.3 Evaluation on Color Frame Reconstruction.
To further demonstrate the ability of our method for color image reconstruction, we use the CED dataset, which utilizes the DAVIS 346c camera to capture color frames and events. This dataset is essential for demonstrating the color reconstruction capability of the proposed method. As shown in Fig. 4, Semantic-E2VID produces more visually compelling reconstructions, particularly in scenes containing colored objects and fast-moving dynamics. The reconstructed images of Semantic-E2VID have higher contrast and clear edges, which proves the effectiveness and robustness of our proposed method.
@@
4.2.4 Evaluation on Inference Speed.
In addition to reconstruction accuracy, inference efficiency is a critical factor for E2V reconstruction, especially in real-time and high-speed vision applications. To evaluate the tradeoff between reconstruction performance and inference speed, we compare Semantic-E2VID with representative E2V methods in terms of both accuracy and frames per second (FPS) on two datasets. Fig. 5 visualizes this tradeoff by jointly plotting reconstruction quality and computational efficiency.
Specifically, the horizontal axis represents the reconstruction error measured by MSE (lower is better), while the vertical axis denotes structural similarity measured by SSIM (higher is better). Inference speed FPS is encoded by the area of each marker, where larger circles indicate faster processing. Methods located toward the upper-left region therefore achieve a more favorable balance between reconstruction accuracy and efficiency.
As shown in Fig. 5, Semantic-E2VID consistently occupies the most desirable region across both ECD and MVSEC datasets. Compared to prior approaches, our method achieves significantly lower reconstruction error and higher structural similarity, while maintaining competitive inference speed. Despite incorporating additional semantic modeling components, Semantic-E2VID does not sacrifice efficiency and remains suitable for practical deployment. These results demonstrate that explicitly modeling semantic information improves reconstruction quality without introducing prohibitive computational overhead, leading to a superior accuracy–efficiency tradeoff.
Figure 5: Inference speed–performance tradeoff for event-to-video reconstruction on two benchmarks: (a) ECD and (b) MVSEC. Each marker corresponds to an E2V method, where the horizontal axis denotes MSE (lower is better, leftward) and the vertical axis denotes SSIM (higher is better, upward). The area of each circle is proportional to the inference speed (FPS), indicating computational efficiency, while different colors represent different methods. Methods located toward the upper-left achieve better reconstruction quality, and larger circles indicate faster inference. Overall, Semantic-E2VID consistently occupies the most favorable region, achieving lower reconstruction error, higher structural similarity, and a superior accuracy–efficiency tradeoff compared to prior approaches.
4.3 Qualitative Analysis
@@
4.3.1 Overall Qualitative Comparison.
Fig. 6 showcases qualitative reconstruction results of our Semantic-E2VID alongside all other SOTA methods on samples from datasets with referenced GT frames. It can be observed that methods such as E2VID, FireNet+, and SPADE-E2VID struggle with brightness consistency, leading to noticeably degraded visual quality. Our Semantic-E2VID demonstrates superior performance in preserving fine-grained object boundaries. This improvement is attributed to the incorporation of rich semantic information, which enhances the overall fidelity of the reconstruction. Furthermore, the outputs of Semantic-E2VID show a closer resemblance to the ground truth frames, delivering high visual quality while effectively mitigating reconstruction artifacts. Semantic-aware guidance allows our model to reconstruct high-level scene understanding, including object shapes and spatial relationships, yielding more realistic and interpretable video frames.
4.3.2 Qualitative Analysis on Reconstruction Quality of Semantic Regions
To further demonstrate the effectiveness of our method in reconstructing semantically meaningful regions, we provide qualitative comparisons with state-of-the-art methods, including E2VID [32], E2VID+ [36], FireNet [33], FireNet+ [36], and ours. As shown in Fig. 7, our method produces more accurate and sharper reconstructions in regions with rich semantic content (e.g., pedestrians, vehicles, or structured indoor objects).
@@ 4.3.2 Qualitative Analysis on Reconstruction Quality of Semantic Regions
4.3.3 Qualitative Analysis on Downstream Segmentation Task
@@
4.3.4 Qualitative Analysis on Semantic Feature Learning
To better understand how our proposed semantic-guided learning module benefits the reconstruction process, we visualize the intermediate feature maps, $\mathcal{F}_{e}^{\prime}$, and final reconstructions from our method, and compare the visualization of the intermediate feature maps $\mathcal{F}_{e}$, and final reconstructions from the baseline model in Fig. 9. We apply colormaps to visualize the feature activations, where warmer colors (e.g., red and yellow) indicate higher activation intensities, and cooler colors (e.g., blue) represent lower activations.